# It’s TIME: Towards the Next Generation of Time Series Forecasting Benchmarks

### v1

@@ Appendix E Full Results
US Term Structure MASE (Raw) 1.458 1.487 1.602 1.542 3.484 1.487 1.563 1.510 1.506 1.586 1.712 1.542
CRPS (Raw) 0.210 0.209 0.224 0.215 0.406 0.208 0.221 0.218 0.214 0.227 0.212 0.218
MASE (Norm.) 0.837 0.854 0.920 0.886 2.001 0.854 0.897 0.867 0.865 0.911 0.983 0.886
CRPS (Norm.) 0.846 0.840 0.903 0.865 1.634 0.837 0.889 0.876 0.863 0.912 0.853 0.876
MASE (Rank) 1 2 10 6 13 3 8 5 4 9 11 7
CRPS (Rank) 3 2 10 6 13 1 9 8 5 11 4 7
Oil Price MASE (Raw) 1.307 1.298 1.343 1.383 2.290 1.312 1.319 1.293 1.344 1.351 1.662 1.476
CRPS (Raw) 0.045 0.045 0.046 0.047 0.088 0.045 0.045 0.046 0.046 0.046 0.057 0.050
MASE (Norm.) 0.871 0.865 0.895 0.921 1.526 0.875 0.879 0.862 0.896 0.900 1.107 0.984
CRPS (Norm.) 0.836 0.836 0.862 0.882 1.651 0.840 0.849 0.869 0.854 0.863 1.070 0.939
MASE (Rank) 3 2 6 9 13 4 5 1 7 8 12 10
CRPS (Rank) 2 1 6 9 13 3 4 8 5 7 12 10
Job Claims MASE (Raw) 2.732 3.153 3.381 2.799 2.978 3.254 2.392 5.572 2.899 4.043 2.641 3.432
CRPS (Raw) 0.018 0.018 0.021 0.017 0.018 0.021 0.015 0.032 0.018 0.022 0.017 0.021
MASE (Norm.) 0.960 1.108 1.188 0.984 1.047 1.144 0.841 1.958 1.019 1.421 0.928 1.206
CRPS (Norm.) 0.934 0.943 1.063 0.855 0.948 1.059 0.782 1.625 0.921 1.127 0.871 1.074
MASE (Rank) 3 8 10 4 7 9 1 13 6 12 2 11
CRPS (Rank) 5 6 10 2 7 9 1 13 4 12 3 11
Uncertainty-1M MASE (Raw) 0.376 0.351 0.339 0.374 0.463 0.436 0.382 0.422 0.448 0.417 0.420 0.551
CRPS (Raw) 0.031 0.027 0.028 0.029 0.040 0.033 0.028 0.035 0.038 0.033 0.032 0.043
MASE (Norm.) 0.355 0.331 0.321 0.354 0.438 0.412 0.361 0.399 0.424 0.394 0.397 0.520
CRPS (Norm.) 0.392 0.343 0.356 0.363 0.505 0.415 0.356 0.443 0.473 0.414 0.402 0.546
MASE (Rank) 4 2 1 3 11 9 5 8 10 6 7 12
CRPS (Rank) 5 1 2 4 11 8 3 9 10 7 6 12
Housing Inventory MASE (Raw) 0.428 0.529 0.512 0.526 0.469 0.579 0.412 1.035 0.566 0.527 0.543 0.496
CRPS (Raw) 0.043 0.053 0.047 0.044 0.045 0.058 0.042 0.100 0.052 0.054 0.049 0.055
MASE (Norm.) 0.595 0.736 0.712 0.732 0.652 0.806 0.573 1.440 0.787 0.733 0.756 0.690
CRPS (Norm.) 0.644 0.792 0.701 0.663 0.671 0.869 0.635 1.499 0.787 0.807 0.733 0.821
MASE (Rank) 2 8 5 6 3 11 1 13 10 7 9 4
CRPS (Rank) 2 8 5 3 4 11 1 13 7 9 6 10
JOLTS MASE (Raw) 1.017 0.906 1.244 1.041 1.205 1.040 1.071 1.546 1.109 1.056 1.501 1.480
CRPS (Raw) 0.063 0.058 0.075 0.065 0.076 0.064 0.068 0.094 0.069 0.065 0.090 0.093
MASE (Norm.) 0.622 0.555 0.761 0.637 0.738 0.637 0.656 0.946 0.679 0.647 0.919 0.906
CRPS (Norm.) 0.628 0.572 0.742 0.642 0.750 0.640 0.672 0.931 0.682 0.645 0.897 0.924
MASE (Rank) 2 1 9 4 8 3 6 12 7 5 11 10
CRPS (Rank) 2 1 8 4 9 3 6 12 7 5 10 11
US Labor MASE (Raw) 0.779 1.014 1.056 0.811 0.908 0.966 0.772 1.452 0.833 0.965 0.903 0.829
CRPS (Raw) 0.042 0.057 0.060 0.046 0.053 0.049 0.044 0.062 0.045 0.056 0.055 0.044
MASE (Norm.) 0.436 0.568 0.592 0.454 0.509 0.541 0.433 0.813 0.466 0.540 0.506 0.464
CRPS (Norm.) 0.380 0.516 0.543 0.409 0.476 0.437 0.397 0.553 0.400 0.505 0.497 0.391
MASE (Rank) 2 10 11 3 7 9 1 12 5 8 6 4
CRPS (Rank) 1 10 11 5 7 6 3 12 4 9 8 2
Vehicle Supply MASE (Raw) 0.798 0.708 0.884 0.757 1.137 0.694 0.764 0.717 0.710 0.813 0.791 0.973
CRPS (Raw) 0.188 0.167 0.236 0.190 0.327 0.159 0.172 0.157 0.157 0.195 0.174 0.278
MASE (Norm.) 0.719 0.638 0.796 0.682 1.024 0.625 0.688 0.646 0.639 0.732 0.712 0.877
CRPS (Norm.) 0.541 0.481 0.679 0.546 0.941 0.458 0.496 0.452 0.452 0.561 0.501 0.800
MASE (Rank) 8 2 10 5 13 1 6 4 3 9 7 11
CRPS (Rank) 7 4 10 8 12 3 5 1 2 9 6 11
Auto Production-SF MASE (Raw) 0.949 0.609 1.228 0.905 0.617 0.963 1.014 1.403 0.992 1.031 1.968 1.330
CRPS (Raw) 0.029 0.019 0.038 0.028 0.020 0.032 0.031 0.046 0.030 0.032 0.061 0.042
MASE (Norm.) 0.971 0.623 1.257 0.926 0.631 0.985 1.038 1.435 1.015 1.055 2.014 1.361
CRPS (Norm.) 0.898 0.599 1.180 0.882 0.621 0.996 0.987 1.451 0.957 1 1.922 1.307
MASE (Rank) 4 1 10 3 2 5 8 12 7 9 13 11
CRPS (Rank) 4 1 10 3 2 7 6 12 5 9 13 11
Commodity Production MASE (Raw) 0.914 0.905 0.928 0.950 0.914 0.912 0.949 1.205 0.794 0.939 0.960 1.027
CRPS (Raw) 0.124 0.124 0.123 0.128 0.124 0.128 0.128 0.151 0.109 0.124 0.128 0.138
MASE (Norm.) 0.734 0.726 0.744 0.763 0.734 0.732 0.761 0.967 0.637 0.753 0.771 0.824
CRPS (Norm.) 0.698 0.695 0.693 0.717 0.699 0.720 0.721 0.851 0.613 0.699 0.720 0.776
@@
CRPS (Rank) 4 3 2 7 6 8 10 12 1 5 9 11
Commodity Import MASE (Raw) 1.567 1.475 1.554 1.453 1.507 1.484 1.499 1.813 1.678 1.534 1.642 1.751
CRPS (Raw) 0.214 0.237 0.213 0.190 0.223 0.195 0.201 0.199 0.185 0.214 0.221 0.197
MASE (Norm.) 0.839 0.790 0.832 0.778 0.807 0.795 0.803 0.971 0.899 0.822 0.880 0.938
CRPS (Norm.) 0.643 0.712 0.641 0.571 0.672 0.587 0.603 0.598 0.555 0.644 0.663 0.591
@@
CRPS (Rank) 8 12 7 2 11 3 6 5 1 9 10 4
WUI-Global MASE (Raw) 0.961 0.946 0.976 0.971 1.005 0.970 0.975 1.022 0.971 0.992 0.959 0.963
CRPS (Raw) 0.513 0.534 0.533 0.531 0.541 0.541 0.553 0.573 0.564 0.547 0.525 0.513
MASE (Norm.) 0.718 0.706 0.729 0.725 0.751 0.725 0.728 0.763 0.725 0.741 0.716 0.719
CRPS (Norm.) 0.623 0.647 0.646 0.644 0.656 0.656 0.671 0.695 0.683 0.664 0.636 0.623
@@
CRPS (Rank) 2 6 5 4 8 7 10 12 11 9 3 1
Global Price MASE (Raw) 1.416 1.430 1.512 1.478 1.555 1.495 1.412 1.748 1.497 1.567 1.464 1.482
CRPS (Raw) 0.132 0.133 0.141 0.140 0.144 0.145 0.131 0.172 0.148 0.146 0.139 0.136
MASE (Norm.) 0.777 0.784 0.829 0.810 0.853 0.820 0.774 0.958 0.821 0.859 0.803 0.812
CRPS (Norm.) 0.735 0.744 0.788 0.779 0.805 0.809 0.732 0.957 0.825 0.813 0.772 0.759
@@
CRPS (Rank) 2 3 7 6 8 9 1 12 11 10 5 4
Vehicle Sales MASE (Raw) 0.861 0.835 0.937 0.759 1.311 0.791 0.750 0.903 0.882 0.773 0.786 0.865
CRPS (Raw) 0.070 0.066 0.073 0.065 0.122 0.067 0.064 0.072 0.068 0.065 0.070 0.075
MASE (Norm.) 0.611 0.592 0.664 0.538 0.930 0.561 0.532 0.640 0.626 0.548 0.557 0.614
CRPS (Norm.) 0.610 0.571 0.640 0.570 1.062 0.581 0.559 0.630 0.594 0.566 0.612 0.652
@@
CRPS (Rank) 7 4 10 3 13 5 1 9 6 2 8 11
Online Retail II MASE (Raw) 0.590 0.609 0.652 0.630 0.733 0.639 0.700 0.765 0.643 0.620 0.671 0.653
CRPS (Raw) 0.234 0.242 0.255 0.247 0.272 0.254 0.274 0.302 0.253 0.248 0.268 0.258
MASE (Norm.) 0.276 0.284 0.305 0.294 0.343 0.298 0.327 0.357 0.300 0.290 0.313 0.305
CRPS (Norm.) 0.234 0.242 0.255 0.247 0.272 0.254 0.274 0.302 0.253 0.248 0.268 0.258
MASE (Rank) 1 2 7 4 11 5 10 12 6 3 9 8
CRPS (Rank) 1 2 7 3 10 6 11 12 5 4 9 8
Supply Chain-Customer MASE (Raw) 0.581 0.554 0.614 0.598 0.636 0.615 0.669 0.666 0.613 0.613 0.661 0.661
CRPS (Raw) 0.275 0.259 0.293 0.276 0.300 0.284 0.311 0.326 0.283 0.292 0.304 0.308
MASE (Norm.) 0.343 0.327 0.363 0.353 0.375 0.363 0.395 0.393 0.362 0.362 0.390 0.390
CRPS (Norm.) 0.258 0.242 0.274 0.258 0.281 0.265 0.291 0.305 0.265 0.273 0.284 0.288
MASE (Rank) 2 1 6 3 8 7 12 11 4 5 9 10
@@
Supply Chain-Location MASE (Raw) 0.627 0.605 0.650 0.649 0.668 0.659 0.699 0.705 0.649 0.650 0.683 0.694
CRPS (Raw) 0.289 0.271 0.305 0.290 0.309 0.295 0.319 0.337 0.291 0.304 0.309 0.318
MASE (Norm.) 0.393 0.379 0.407 0.406 0.418 0.413 0.438 0.441 0.406 0.407 0.428 0.434
CRPS (Norm.) 0.273 0.257 0.288 0.274 0.292 0.279 0.302 0.319 0.275 0.287 0.293 0.300
MASE (Rank) 2 1 5 4 8 7 11 12 3 6 9 10
CRPS (Rank) 2 1 7 3 8 5 11 12 4 6 9 10
Azure2019-D MASE (Raw) 0.731 0.716 0.834 0.758 0.899 0.763 0.741 0.776 0.767 0.807 0.869 0.817
@@
MASE (Norm.) 0.668 0.654 0.762 0.693 0.822 0.697 0.677 0.709 0.701 0.738 0.795 0.747
CRPS (Norm.) 0.492 0.473 0.662 0.518 0.630 0.528 0.494 0.568 0.550 0.561 0.599 0.621
@@
MASE (Norm.) 0.690 0.669 0.779 0.718 0.784 0.719 0.694 0.753 0.773 0.780 0.906 1.114
CRPS (Norm.) 0.565 0.538 0.673 0.584 0.637 0.585 0.563 0.639 0.641 0.640 0.658 0.805
MASE (Rank) 2 1 8 4 10 5 3 6 7 9 11 13
@@
MASE (Norm.) 0.602 0.574 0.662 0.588 0.737 0.607 0.594 0.680 0.626 0.671 0.781 0.740
CRPS (Norm.) 0.283 0.257 0.328 0.260 0.349 0.264 0.269 0.295 0.286 0.314 0.331 0.320
MASE (Rank) 4 1 7 2 10 5 3 9 6 8 12 11
@@
Smart Manufacturing MASE (Raw) 0.713 0.712 0.713 0.714 0.716 0.719 0.720 0.734 0.717 0.715 0.728 0.717
CRPS (Raw) 0.167 0.166 0.175 0.167 0.167 0.167 0.169 0.188 0.178 0.173 0.172 0.180
MASE (Norm.) 0.685 0.685 0.685 0.686 0.688 0.691 0.692 0.706 0.689 0.687 0.700 0.689
CRPS (Norm.) 0.610 0.610 0.641 0.613 0.612 0.614 0.618 0.689 0.653 0.632 0.629 0.657
MASE (Rank) 2 1.667 3.333 3.667 6.667 9 9.333 11.667 6.667 5.333 11.333 7.333
CRPS (Rank) 3 1.667 7.333 4 2.667 5.333 7 11.333 9.333 7 8.667 10.667
MetroPT-3 MASE (Raw) 0.625 0.629 0.666 0.665 0.673 0.645 0.645 0.669 0.690 0.689 0.673 0.711
CRPS (Raw) 0.256 0.255 0.288 0.253 0.264 0.256 0.262 0.313 0.300 0.280 0.261 0.324
MASE (Norm.) 0.740 0.745 0.788 0.787 0.797 0.763 0.764 0.792 0.817 0.815 0.797 0.842
CRPS (Norm.) 0.677 0.674 0.762 0.670 0.698 0.679 0.693 0.829 0.790 0.742 0.691 0.854
MASE (Rank) 1.667 1.667 7 7 8 3.667 4.333 7.333 9.667 9.333 7 11.333
CRPS (Rank) 2 2.333 8.333 2.667 6.667 3.667 6 11 8.667 9 6.333 11.333

### v2

@@ Appendix E Full Results
US Term Structure MASE (raw) 1.458 1.460 1.602 1.542 3.484 1.479 1.563 1.510 1.506 1.535 1.712 1.542
CRPS (raw) 0.210 0.209 0.224 0.215 0.406 0.210 0.221 0.218 0.214 0.221 0.212 0.218
MASE (norm.) 0.837 0.839 0.920 0.886 2.001 0.849 0.897 0.867 0.865 0.882 0.983 0.886
CRPS (norm.) 0.846 0.842 0.903 0.865 1.634 0.844 0.889 0.876 0.863 0.891 0.853 0.876
MASE (rank) 1 2 10 7 13 3 9 5 4 6 11 8
CRPS (rank) 3 1 11 6 13 2 9 8 5 10 4 7
Oil Price MASE (raw) 1.307 1.300 1.343 1.383 2.290 1.333 1.319 1.293 1.344 1.376 1.662 1.476
CRPS (raw) 0.045 0.045 0.046 0.047 0.088 0.046 0.045 0.046 0.046 0.046 0.057 0.050
MASE (norm.) 0.871 0.866 0.895 0.921 1.526 0.889 0.879 0.862 0.896 0.917 1.107 0.984
CRPS (norm.) 0.836 0.836 0.862 0.882 1.651 0.859 0.849 0.869 0.854 0.863 1.070 0.939
MASE (rank) 3 2 6 9 13 5 4 1 7 8 12 10
CRPS (rank) 2 1 6 9 13 5 3 8 4 7 12 10
Job Claims MASE (raw) 2.732 3.144 3.381 2.799 2.978 2.431 2.392 5.572 2.899 4.147 2.641 3.432
CRPS (raw) 0.018 0.018 0.021 0.017 0.018 0.016 0.015 0.032 0.018 0.022 0.017 0.021
MASE (norm.) 0.960 1.105 1.188 0.984 1.047 0.854 0.841 1.958 1.019 1.458 0.928 1.206
CRPS (norm.) 0.934 0.947 1.063 0.855 0.948 0.848 0.782 1.625 0.921 1.145 0.871 1.074
MASE (rank) 4 9 10 5 8 2 1 13 7 12 3 11
CRPS (rank) 6 7 10 3 8 2 1 13 5 12 4 11
Uncertainty-1M MASE (raw) 0.376 0.348 0.339 0.374 0.463 0.371 0.382 0.422 0.448 0.415 0.420 0.551
CRPS (raw) 0.031 0.027 0.028 0.029 0.040 0.027 0.028 0.035 0.038 0.033 0.032 0.043
MASE (norm.) 0.355 0.329 0.321 0.354 0.438 0.351 0.361 0.399 0.424 0.392 0.397 0.520
CRPS (norm.) 0.392 0.340 0.356 0.363 0.505 0.342 0.356 0.443 0.473 0.414 0.402 0.546
MASE (rank) 5 2 1 4 11 3 6 9 10 7 8 12
CRPS (rank) 6 1 3 5 11 2 4 9 10 8 7 12
Housing Inventory MASE (raw) 0.428 0.530 0.512 0.526 0.469 0.483 0.412 1.035 0.566 0.528 0.543 0.496
CRPS (raw) 0.043 0.053 0.047 0.044 0.045 0.044 0.042 0.100 0.052 0.054 0.049 0.055
MASE (norm.) 0.595 0.737 0.712 0.732 0.652 0.672 0.573 1.440 0.787 0.735 0.756 0.690
CRPS (norm.) 0.644 0.791 0.701 0.663 0.671 0.658 0.635 1.499 0.787 0.806 0.733 0.821
MASE (rank) 2 9 6 7 3 4 1 13 11 8 10 5
CRPS (rank) 2 9 6 4 5 3 1 13 8 10 7 11
JOLTS MASE (raw) 1.017 0.904 1.244 1.041 1.205 0.995 1.071 1.546 1.109 1.055 1.501 1.480
CRPS (raw) 0.063 0.058 0.075 0.065 0.076 0.062 0.068 0.094 0.069 0.065 0.090 0.093
MASE (norm.) 0.622 0.554 0.761 0.637 0.738 0.609 0.656 0.946 0.679 0.646 0.919 0.906
CRPS (norm.) 0.628 0.572 0.742 0.642 0.750 0.614 0.672 0.931 0.682 0.645 0.897 0.924
MASE (rank) 3 1 9 4 8 2 6 12 7 5 11 10
CRPS (rank) 3 1 8 4 9 2 6 12 7 5 10 11
US Labor MASE (raw) 0.779 1.001 1.056 0.811 0.908 0.790 0.772 1.452 0.833 0.946 0.903 0.829
CRPS (raw) 0.042 0.058 0.060 0.046 0.053 0.044 0.044 0.062 0.045 0.056 0.055 0.044
MASE (norm.) 0.436 0.561 0.592 0.454 0.509 0.443 0.433 0.813 0.466 0.530 0.506 0.464
CRPS (norm.) 0.380 0.517 0.543 0.409 0.476 0.393 0.397 0.553 0.400 0.504 0.497 0.391
MASE (rank) 2 10 11 4 8 3 1 12 6 9 7 5
CRPS (rank) 1 10 11 6 7 3 4 12 5 9 8 2
Vehicle Supply MASE (raw) 0.798 0.709 0.884 0.757 1.137 0.792 0.764 0.717 0.710 0.811 0.791 0.973
CRPS (raw) 0.188 0.167 0.236 0.190 0.327 0.193 0.172 0.157 0.157 0.195 0.174 0.278
MASE (norm.) 0.719 0.639 0.796 0.682 1.024 0.714 0.688 0.646 0.639 0.730 0.712 0.877
CRPS (norm.) 0.541 0.482 0.679 0.546 0.941 0.556 0.496 0.452 0.452 0.560 0.501 0.800
MASE (rank) 8 1 10 4 13 7 5 3 2 9 6 11
CRPS (rank) 6 3 10 7 12 8 4 1 2 9 5 11
Auto Production-SF MASE (raw) 0.949 0.604 1.228 0.905 0.617 0.746 1.014 1.403 0.992 1.032 1.968 1.330
CRPS (raw) 0.029 0.019 0.038 0.028 0.020 0.023 0.031 0.046 0.030 0.032 0.061 0.042
MASE (norm.) 0.971 0.618 1.257 0.926 0.631 0.763 1.038 1.435 1.015 1.056 2.014 1.361
CRPS (norm.) 0.898 0.597 1.180 0.882 0.621 0.720 0.987 1.451 0.957 1 1.922 1.307
MASE (rank) 5 1 10 4 2 3 8 12 7 9 13 11
CRPS (rank) 5 1 10 4 2 3 7 12 6 9 13 11
Commodity Production MASE (raw) 0.914 0.905 0.928 0.950 0.914 0.914 0.949 1.205 0.794 0.939 0.960 1.027
CRPS (raw) 0.124 0.124 0.123 0.128 0.124 0.124 0.128 0.151 0.109 0.124 0.128 0.138
MASE (norm.) 0.734 0.726 0.744 0.763 0.734 0.733 0.761 0.967 0.637 0.753 0.771 0.824
CRPS (norm.) 0.698 0.695 0.693 0.717 0.699 0.695 0.721 0.851 0.613 0.699 0.720 0.776
@@
CRPS (rank) 5 3 2 8 7 4 10 12 1 6 9 11
Commodity Import MASE (raw) 1.567 1.459 1.554 1.453 1.507 1.498 1.499 1.813 1.678 1.532 1.642 1.751
CRPS (raw) 0.214 0.238 0.213 0.190 0.223 0.226 0.201 0.199 0.185 0.214 0.221 0.197
MASE (norm.) 0.839 0.782 0.832 0.778 0.807 0.802 0.803 0.971 0.899 0.821 0.880 0.938
CRPS (norm.) 0.643 0.714 0.641 0.571 0.672 0.678 0.603 0.598 0.555 0.644 0.663 0.591
@@
CRPS (rank) 7 12 6 2 10 11 5 4 1 8 9 3
WUI-Global MASE (raw) 0.961 0.949 0.976 0.971 1.005 0.966 0.975 1.022 0.971 0.998 0.959 0.963
CRPS (raw) 0.513 0.534 0.533 0.531 0.541 0.521 0.553 0.573 0.564 0.550 0.525 0.513
MASE (norm.) 0.718 0.708 0.729 0.725 0.751 0.721 0.728 0.763 0.725 0.745 0.716 0.719
CRPS (norm.) 0.623 0.647 0.646 0.644 0.656 0.631 0.671 0.695 0.683 0.667 0.636 0.623
@@
CRPS (rank) 2 7 6 5 8 3 10 12 11 9 4 1
Global Price MASE (raw) 1.416 1.427 1.512 1.478 1.555 1.496 1.412 1.748 1.497 1.568 1.464 1.482
CRPS (raw) 0.132 0.133 0.141 0.140 0.144 0.144 0.131 0.172 0.148 0.146 0.139 0.136
MASE (norm.) 0.777 0.783 0.829 0.810 0.853 0.820 0.774 0.958 0.821 0.860 0.803 0.812
CRPS (norm.) 0.735 0.742 0.788 0.779 0.805 0.800 0.732 0.957 0.825 0.815 0.772 0.759
@@
CRPS (rank) 2 3 7 6 9 8 1 12 11 10 5 4
Vehicle Sales MASE (raw) 0.861 0.835 0.937 0.759 1.311 0.803 0.750 0.903 0.882 0.772 0.786 0.865
CRPS (raw) 0.070 0.065 0.073 0.065 0.122 0.066 0.064 0.072 0.068 0.065 0.070 0.075
MASE (norm.) 0.611 0.592 0.664 0.538 0.930 0.570 0.532 0.640 0.626 0.548 0.557 0.614
CRPS (norm.) 0.610 0.570 0.640 0.570 1.062 0.573 0.559 0.630 0.594 0.566 0.612 0.652
@@
CRPS (rank) 7 3 10 4 13 5 1 9 6 2 8 11
Online Retail II MASE (raw) 0.590 0.609 0.652 0.630 0.733 0.609 0.700 0.765 0.643 0.620 0.671 0.653
CRPS (raw) 0.234 0.242 0.255 0.247 0.272 0.244 0.274 0.302 0.253 0.248 0.268 0.258
MASE (norm.) 0.276 0.285 0.305 0.294 0.343 0.284 0.327 0.357 0.300 0.289 0.313 0.305
CRPS (norm.) 0.234 0.242 0.255 0.247 0.272 0.245 0.274 0.302 0.253 0.248 0.268 0.258
MASE (rank) 1 3 7 5 11 2 10 12 6 4 9 8
CRPS (rank) 1 2 7 4 10 3 11 12 6 5 9 8
Supply Chain-Customer MASE (raw) 0.581 0.553 0.614 0.598 0.636 0.604 0.669 0.666 0.613 0.613 0.661 0.661
CRPS (raw) 0.275 0.259 0.293 0.276 0.300 0.281 0.311 0.326 0.283 0.292 0.304 0.308
MASE (norm.) 0.343 0.327 0.363 0.353 0.375 0.356 0.395 0.393 0.362 0.362 0.390 0.390
CRPS (norm.) 0.258 0.242 0.274 0.258 0.281 0.263 0.291 0.305 0.265 0.273 0.284 0.288
MASE (rank) 2 1 7 3 8 4 12 11 5 6 9 10
CRPS (rank) 2 1 7 3 8 4 11 12 5 6 9 10
Supply Chain-Location MASE (raw) 0.627 0.605 0.650 0.649 0.668 0.649 0.699 0.705 0.649 0.650 0.683 0.694
CRPS (raw) 0.289 0.271 0.305 0.290 0.309 0.292 0.319 0.337 0.291 0.304 0.309 0.318
MASE (norm.) 0.393 0.379 0.407 0.406 0.418 0.406 0.438 0.441 0.406 0.407 0.428 0.434
CRPS (norm.) 0.273 0.257 0.288 0.274 0.292 0.276 0.302 0.319 0.275 0.287 0.293 0.300
MASE (rank) 2 1 6 5 8 3 11 12 4 7 9 10
@@
Azure2019-D MASE (raw) 0.731 0.718 0.834 0.758 0.899 0.763 0.741 0.776 0.767 0.807 0.869 0.817
@@
MASE (norm.) 0.668 0.656 0.762 0.693 0.822 0.697 0.677 0.709 0.701 0.738 0.795 0.747
CRPS (norm.) 0.492 0.474 0.662 0.518 0.630 0.528 0.494 0.568 0.550 0.561 0.599 0.621
@@
MASE (norm.) 0.690 0.667 0.779 0.718 0.784 0.719 0.694 0.753 0.773 0.779 0.906 1.114
CRPS (norm.) 0.565 0.537 0.673 0.584 0.637 0.585 0.563 0.639 0.641 0.640 0.658 0.805
MASE (rank) 2 1 9 4 10 5 3 6 7 8 11 13
@@
MASE (norm.) 0.602 0.574 0.662 0.588 0.737 0.581 0.594 0.680 0.626 0.671 0.781 0.740
CRPS (norm.) 0.283 0.257 0.328 0.260 0.349 0.263 0.269 0.295 0.286 0.314 0.331 0.320
MASE (rank) 5 1 7 3 10 2 4 9 6 8 12 11
@@
Smart Manufacturing MASE (raw) 0.713 0.711 0.713 0.714 0.716 0.716 0.720 0.734 0.717 0.715 0.728 0.717
CRPS (raw) 0.167 0.166 0.175 0.167 0.167 0.167 0.169 0.188 0.178 0.170 0.172 0.180
MASE (norm.) 0.685 0.684 0.685 0.686 0.688 0.688 0.692 0.706 0.689 0.687 0.700 0.689
CRPS (norm.) 0.610 0.608 0.641 0.613 0.612 0.611 0.618 0.689 0.653 0.623 0.629 0.657
MASE (rank) 2.333 1 3.333 3.667 7 7 9.667 11.667 6.333 6.333 11.333 8.333
CRPS (rank) 3.667 1 7.333 4.333 3.667 3.667 7 11.333 9.667 7 8.667 10.667
MetroPT-3 MASE (raw) 0.625 0.610 0.666 0.665 0.673 0.616 0.645 0.669 0.690 0.641 0.673 0.711
CRPS (raw) 0.256 0.244 0.288 0.253 0.264 0.246 0.262 0.313 0.300 0.264 0.261 0.324
MASE (norm.) 0.740 0.722 0.788 0.787 0.797 0.729 0.764 0.792 0.817 0.760 0.797 0.842
CRPS (norm.) 0.677 0.647 0.762 0.670 0.698 0.652 0.693 0.829 0.790 0.698 0.691 0.854
MASE (rank) 2.667 1.333 8 8 8.667 2.667 5.333 8.333 10 4.667 7 11.333
CRPS (rank) 3.667 1 8.667 4 7 2 6.667 11 9 6.667 6.667 11.667